# labelinglabel

### v1

@@ Abstract
Large Language Models (LLMs) are increasingly used as evaluators in natural language processing tasks, a paradigm known as LLM-as-a-Judge (LaaJ). While effective in general domains, LaaJs pose significant challenges in domain-specific contexts, where annotated data is scarce and expert evaluation is costly. In such cases, meta-evaluation is often performed using metrics that have not been validated for the specific domain in which they are applied. As a result, it becomes difficult to determine which metrics effectively identify LaaJ quality, and further, what threshold indicates sufficient evaluator performance.
In this work, we introduce LaaJMeter, a simulation-based framework for controlled meta-evaluation of LaaJs. LaaJMeter enables engineers to generate synthetic data representing virtual models and judges, allowing systematic analysis of evaluation metrics under realistic conditions. This helps practitioners validate and refine LaaJs for specific evaluation tasks: they can test whether their metrics correctly distinguish between better and worse (virtual) LaaJs, and estimate appropriate thresholds for evaluator adequacy.
@@ 1 Introduction
Large Language Models (LLMs) have dramatically expanded the capabilities of Natural Language Processing (NLP), making a wide range of tasks such as translation, summarization, question answering, and code generation not only feasible but scalable and accessible. Their ability to generalize across domains and adapt to diverse linguistic inputs has enabled the deployment of NLP tools in settings where traditional rule-based or supervised approaches were previously impractical.
@@
As NLP systems become more pervasive, the need to evaluate their outputs reliably and efficiently becomes increasingly critical. Importantly, evaluating an NLP tool, whether AI-based or not, is itself an NLP task. It involves interpreting textual outputs, assessing their correctness, relevance, or fluency, and making nuanced judgments that often depend on domain-specific knowledge. This evaluative layer is essential for validating system behavior, guiding model development, and ensuring deployment safety [14].
LLMs are now widely used as evaluators, a paradigm known as LLM as a Judge (LaaJ), where the model is tasked with scoring, ranking, or otherwise assessing the outputs of other systems [17]. This approach has proven effective and is increasingly adopted in both research and production environments [5]. LaaJs are valued primarily for their scalability: they can provide context-aware evaluations across large datasets without the need for manual annotation. Beyond standalone evaluation, LaaJs are also integrated into AI-agent architectures, where they assess the outputs of other components and influence downstream decisions. In such systems, the judgment of a LaaJ may directly affect the agent’s actions, making the reliability of these judgments critical [16].
However, the growing reliance on LaaJs introduces a critical question: who evaluates the evaluator? This challenge, known as meta-evaluation, concerns the assessment of the judgments produced by an LLM acting as a judge [5]. A robust meta-evaluation framework must determine whether these judgments are reliable, consistent, and meaningful across diverse tasks and inputs. A tempting but flawed solution is to use another LLM for this purpose. Yet, doing so merely shifts the problem up a level, creating a recursive loop with no independent ground truth. Without an external, model-agnostic mechanism for evaluation, the validity of the entire process is called into question. Thus, principled and reproducible approaches to meta-evaluation must break this dependency cycle and avoid relying on LLMs themselves.
@@
One attempt to perform meta-evaluation involves using text-similarity metrics such as BLEU, ROUGE, METEOR, and BERTScore to compare LaaJ outputs against reference judgments [13]. While these metrics offer a convenient and automated way to assess textual alignment, they fall short in capturing the nuanced, context-dependent nature of evaluative tasks. For instance, BLEU and ROUGE rely heavily on n-gram overlap, which may penalize valid paraphrases or reward superficial similarity. BERTScore improves upon this by leveraging contextual embeddings, yet it still lacks the ability to assess reasoning quality, factual correctness, or domain-specific relevance. These limitations are particularly pronounced in tasks where multiple valid judgments exist or where subtle distinctions matter. For example, an instruction to write code in a programming language can yield multiple correct solutions, none of which may be textually similar. Therefore, for reliable meta-evaluation, comparison against expert human annotations remains the gold standard [14]. In practice, metrics such as agreement with expert rankings (e.g., via Kendall-$\tau$ or Spearman-$\rho$) and accuracy against gold-standard labels are commonly used to benchmark LaaJ performance against human judgment. Moreover, off-the-shelf benchmarks exist for many common tasks such as summarization (e.g., CNN/DailyMail), question answering (e.g., SQuAD), and code generation (e.g., HumanEval), providing standardized datasets and annotations for principled evaluation [1, 4].
@@
While off-the-shelf benchmarks provide robust infrastructure for evaluating LaaJs on general tasks such as summarization, question answering, and code generation, they are often ill-suited for domain-specific variants of these tasks. For example, generating code in a legacy programming language, summarizing legal documents, or answering questions in a specialized scientific field may require domain expertise and context that standard benchmarks do not capture [11, 8]. These specialized tasks frequently suffer from data scarcity and limited access to expert annotations, due to the rarity and cost of qualified evaluators. This challenge is especially pronounced in low-budget projects, where the resources needed to curate domain-specific datasets or involve experts are not readily available. As a result, performing principled meta-evaluation in these contexts remains a significant hurdle, demanding creative solutions that can compensate for the lack of standardized resources.
@@
In the absence of standardized benchmarks and expert annotations, meta-evaluation in domain-specific contexts often relies on creative, task-specific strategies. One illustrative approach is to use outputs of known or controlled quality, such as responses from models with varying capabilities, and verify that the LaaJ assigns higher scores to the better outputs. However, due to the specificity of the domain, there is often little to no literature guiding which metrics are effective or which experimental setups yield meaningful insights for meta-evaluation [7]. Even when a metric appears to correlate with human judgment (though this correlation is rarely verifiable), determining a threshold that signifies sufficient evaluator quality remains inherently difficult. Without clear standards, practitioners are left to define ad hoc criteria, which may vary widely across applications and compromise reproducibility.
@@
To address this challenge, we propose LaaJMeter, a simulation-based framework for controlled meta-evaluation of LaaJs. LaaJMeter generates synthetic data representing virtual models and virtual LaaJs, enabling systematic analysis of evaluation metrics under realistic conditions. By simulating models of varying quality and LaaJs with different noise and bias profiles, we can test the sensitivity and robustness of metrics without relying on human annotation [10].
@@
We demonstrate the utility of LaaJMeter through a use case in code translation, where an LLM translates legacy code into a modern programming language. In this context, we simulate virtual models and LaaJs, and evaluate several metrics: t-test, Kendall’s $\tau$ correlation, and an ordering experiment. Our results reveal that while some metrics (e.g., t-test) are insufficiently sensitive to LaaJ quality, others (e.g., $\tau$-correlation) provide robust signals even when model distance estimates are imprecise. The ordering experiment, though effective, requires careful interpretation of model distance to ensure meaningful results.
@@
In summary, LaaJMeter offers a principled and extensible framework for evaluating LLMs as judges, particularly in settings where annotated data is scarce or unreliable. It complements existing evaluation paradigms and provides practical guidance for metric selection, thresholding, and simulation-based benchmarking. Our work contributes to the growing body of research on LLM-based evaluation and opens new directions for scalable, data-efficient assessment of model judgment quality.
2 Related Work
Several prior works collectively underscore the importance of principled, adaptable meta-evaluation strategies for LaaJs, when expert data is limited or unavailable.
The paradigm of LLM-as-a-Judge (LaaJ) has gained traction as a scalable alternative to human evaluation, particularly in domains where expert annotations are scarce. While general-purpose benchmarks such as AlpacaEval and Chatbot Arena have been widely adopted, they often fail to capture the nuances of domain-specific tasks. The work [12] addresses this limitation by introducing a data pipeline for constructing domain-specific evaluation sets tailored for LaaJ frameworks. Their method combines manual curation, semi-supervised clustering, and stratified sampling to ensure balanced representation across domains such as law, medicine, and multilingual contexts.
In the legal domain, the work [3] demonstrates how LLMs can be leveraged for term extraction under limited annotated data conditions. Their study compares prompt-engineering and fine-tuning strategies across several models, showing that zero-shot and one-shot capabilities can significantly reduce reliance on expert-labeled data.
The work [2] explores the use of open-source LLMs for text annotation in political science, highlighting the effectiveness of fine-tuning in low-resource settings. Their findings suggest that modest quantities of annotated data can yield competitive performance, offering a practical path for domain-specific evaluation.
From a broader perspective, the work [9] provides a systematic survey of LLM evaluation methodologies, identifying key challenges in reproducibility and reliability. Their recommendations emphasize the need for standardized protocols, especially in specialized domains where evaluation setups vary widely.
Finally, the work [15] proposes a Bayesian evaluation framework to assess LLMs’ capabilities as functional approximators. Their work offers insights into how prior domain knowledge influences model behavior, which is particularly relevant for designing evaluators in domain-specific contexts.
3 Problem Description
@@
In the absence of a comprehensive expert evaluation, one way to assess LaaJ quality is by applying it to outputs of known, varying quality. The evaluated LaaJ is expected to correctly distinguish between different outputs to the same input, by assigning higher scores to better outputs. This form of meta-evaluation enables comparison between different LaaJs based on their ability to rank outputs in accordance with their true quality. See [6] for an example of such an evaluation. However, even when relative comparisons are possible, it remains difficult to define a satisfactory threshold for what constitutes acceptable LaaJ quality.
@@
4 LaaJMeter, a General Description
To address the problem described above, we propose the creation of synthetic data representing virtual models and virtual LaaJs. This data is designed to simulate the behavior of models and LaaJs on a specific task, enabling controlled meta-evaluation. The key design principle is to ensure that the generated data closely resembles a realistic distribution and the behaviors observed in the evaluated NLP task.
4.1 Virtual Points
@@ 4.2 Virtual Models
4.3 Additional Virtual Models
To simulate models of varying quality, we generate additional virtual models by systematically modifying the scores of an existing model. For example, given a model $M$, we can define a new model $M^{\prime}$ such that $M^{\prime}(k)=M(k)+b_{k}$, where $b_{k}$ is sampled from a distribution. If the distribution tends to be positive, $M^{\prime}$ represents a model of higher quality than $M$. Importantly, this approach maintains a pointwise relationship between models, ensuring that $M^{\prime}$ performs better than $M$ in expectation across all points, mimicking the behavior of improved model versions.
4.4 Virtual LaaJs
@@
5 Code Translation Use Case
We now describe a concrete instantiation of the technique outlined above, applied in the context of evaluating an LLM designed to perform code translation from a legacy programming language to a modern one.
5.1 Virtual Models
We began by constructing a virtual base model $M_{0}:\{0,\dots,99\}\rightarrow[0,30]$. This model was created by assigning ground-truth scores to point indices, sampled from the score distribution of a LaaJ we developed for the code translation task. Accordingly, we chose the range $[0,30]$ to match the scoring scale used by that LaaJ. The number of points, 100, is typical to the benchmarks we use for model evaluation.
Following the construction of $M_{0}$, we generated additional virtual models $M_{1},\dots,M_{20}$. Each model $M_{i+1}$ was derived from $M_{i}$ by modifying its scores as follows: for each point $k$, with probability $p$, we set $M_{i+1}(k)=\min\{M_{i}(k)+1,30\}$, incrementing the score unless it exceeds the maximum value of 30. With probability 1p1-p, we set $M_{i+1}(k)=\max\{M_{i}(k)-1,0\}$, decrementing the score unless it falls below 0. The value of $p$ was chosen such that the expected mean score of $M_{i+1}$ is expected to be 0.5 points higher than that of $M_{i}$, accounting for the number of scores at the boundaries (0 and 30).
5.2 LaaJ Simulation
We now describe the construction of virtual LaaJs, informed by our experience evaluating LaaJs in the code modernization task. In practice, benchmarks often contain a mix of straightforward and challenging examples. A competent LaaJ typically performs well on simple cases, while its performance may vary on examples involving specific features, such as uncommon commands, that introduce evaluation difficulty. Some LaaJs may handle certain features effectively but struggle with others.
6 Code Translation Use Case: Metric Simulation and Findings
6.1 $t$-Test
Importantly, the distance between the models also plays a critical role. If the distance is too large, even a poor LaaJ may detect the difference, leading to a significant $p$-value. To explore this, we simulated $t$-tests on virtual LaaJ scores across pairs of virtual models with varying distances. The resulting $p$-values are shown in Table 1.
@@ 6.1 $t$-Test
Table 1: $p$-value results
6.1.1 $t$-Test Result Analysis and Conclusions
As shown in Table 1, all LaaJs yield statistically significant $p$-values for model distances greater than 2. This indicates that even a poor LaaJ, such as $L_{10}$, can detect differences between models, even when the distance is merely notable. Consequently, the $t$-test lacks sensitivity to LaaJ quality and fails to distinguish between LaaJs of varying performance. We conclude that the $t$-test is not suitable for meta-evaluation in this context.
6.2 Kendall-$\tau$ rank correlation
Moreover, we expect the difference in $\tau$-correlation between good and poor LaaJs to be more pronounced when comparing models with small distances. In such cases, subtle differences in inference quality are harder to detect, and only better LaaJs are expected to capture them effectively.
The results of our simulation are presented in Table 2.

### v2

# LaaJMeter: A Framework for LaaJ Evaluation

@@ Abstract
Large Language Models (LLMs) are increasingly used as evaluators in natural language processing tasks, a paradigm known as LLM-as-a-Judge (LaaJ). The analysis of a LaaJ software, commonly refereed to as meta-evaluation, pose significant challenges in domain-specific contexts. In such domains, in contrast to general domains, annotated data is scarce and expert evaluation is costly. As a result, meta-evaluation is often performed using metrics that have not been validated for the specific domain in which they are applied. Therefore, it becomes difficult to determine which metrics effectively identify LaaJ quality, and further, what threshold indicates sufficient evaluator performance.
In this work, we introduce LaaJMeter, a simulation-based framework for controlled meta-evaluation of LaaJs. LaaJMeter enables engineers to generate synthetic data representing virtual models and judges, allowing systematic analysis of evaluation metrics under realistic conditions. This helps practitioners validate LaaJs for specific tasks: they can test whether their metrics correctly distinguish between high and low quality (virtual) LaaJs, and estimate appropriate thresholds for evaluator adequacy.
@@ I Introduction
Large Language Models (LLMs) [NZY21, attention] have dramatically expanded the capabilities of Natural Language Processing (NLP), making a wide range of tasks such as translation, summarization, question answering, and code generation not only feasible but scalable and accessible [MK25, NKQSAUABM25]. Their ability to generalize across domains and adapt to diverse linguistic inputs has enabled the deployment of NLP tools in settings where traditional rule-based or analytical methods were previously impractical.
@@
As NLP systems become increasingly pervasive, the need for reliable and efficient evaluation of their outputs grows more critical. Importantly, evaluating an NLP system, whether AI-based or not, is itself an NLP task. It involves interpreting textual outputs, assessing their correctness, relevance, and fluency, and making nuanced judgments that often depend on domain-specific expertise. This evaluative layer is essential for validating system behavior, guiding model development, and ensuring safe deployment [cambridge2023usereval].
@@
One common evaluation strategy employs text-similarity metrics such as BLEU [BLEU], ROUGE [ROUGE], METEOR [METEOR], and BERTScore [BERTscore], which compare system outputs against reference texts. While these metrics offer a convenient and automated means of assessing textual alignment, they often fail to capture the nuanced, context-dependent nature of evaluative tasks [BDMOS2022, CPL2020, EBSB23, saadany2021metrics]. BLEU and ROUGE rely heavily on n-gram overlap, which can penalize valid paraphrases or reward superficial similarity. BERTScore improves upon this by leveraging contextual embeddings, yet it still lacks the ability to evaluate reasoning quality, factual accuracy, or domain-specific relevance. These limitations are especially pronounced in tasks with multiple valid outputs or where subtle distinctions are critical. For instance, a prompt to generate code in a specific programming language may yield several correct solutions, none of which are textually similar.
A second approach uses off-the-shelf benchmarks that rely on analytic assessments or expert-annotated labels. HumanEval [HumanEval] evaluates code generation using unit tests. GLUE [GLUE] and SQuAD [SQuaD] employ human-annotated data to assess text understanding and reading comprehension, respectively. CNN/DailyMail [HermannKGEKSB15] and WMT [KocmiABBDFFFGGH24] use text similarity metrics like BLEU, ROUGE, and COMMET [ReiSFL20] against gold-standard human data to evaluate bullet summarization and translation, respectively (text similarity metrics are effective for these tasks, as correct outputs are expected to closely match the reference answers). Such benchmarks mainly exist for common NLP tasks.
When annotated data is unavailable, which is often the case for domain-specific tasks, a third approach involves recruiting experts to create dedicated datasets. While this method provides high-quality assessments, it is frequently prohibitively costly. Moreover, when specialized expertise is required, such as for translation or code generation in legacy languages, finding qualified annotators can be challenging.
This paper focuses on a fourth approach. Given that evaluating NLP tools is itself an NLP task, and considering the remarkable success of LLMs in solving such tasks, there is growing interest in using LLMs for evaluation, a paradigm known as LLM-as-a-Judge (LaaJ) [CCL23]. This approach has proven effective and is increasingly adopted in both research and production environments [LIXWZ23, ZCSZWZLLLXZGS23, CWJSX23, HKPA24]. LaaJs are valued primarily for their scalability and cost-efficiency: they can provide context-aware evaluations across large datasets without requiring manual annotation or being limited to benchmark outputs. Furthermore, it avoids the risk of models being trained on the same benchmarks used for evaluation, a concern inherent in the second approach. Beyond standalone evaluation, LaaJs are also integrated into AI-agent architectures, where they assess the outputs of other components and influence downstream decisions. In such systems, the judgment of a LaaJ may directly affect the agent’s actions, making the reliability of these judgments critical.
However, the growing reliance on LaaJs introduces a critical question: who evaluates the evaluator? This challenge, known as meta-evaluation, concerns the assessment of the judgments produced by an LLM acting as a judge [chern2024scaleeval]. A robust meta-evaluation framework must determine whether these judgments are reliable, consistent, and meaningful across diverse tasks and inputs. A tempting but flawed solution is to use another LLM for this purpose. Yet doing so merely shifts the problem up a level, creating a recursive loop with no independent ground truth. Without an external, model-agnostic mechanism for evaluation, the validity of the entire process is called into question. Therefore, principled and reproducible approaches to meta-evaluation must break this dependency cycle and avoid relying on LLMs themselves.
For meta-evaluation, accessible benchmarks and human-annotated data are reused as follows [cambridge2023usereval, GJSTZXLSMLW24]. (1) The evaluated NLP tool is applied on benchmark or annotated data inputs; (2) the LaaJ assesses the tool outputs; (3) measures such as agreement, precision, recall, Cohen’s Kappa, and Spearman’s correlation are used to compare the alignment of LaaJ responses with human labels or benchmark results [BYCLHWYZXLZLH23, LZGSVKC24, TCRVH24].
@@
Consequently, shifting the evaluation task also shifts the underlying challenge: specialized tasks frequently suffer from data scarcity and limited access to expert annotations, due to the rarity and cost of qualified evaluators. This issue is especially pronounced in low-budget projects, where the resources needed to curate domain-specific datasets or engage experts are not readily available. As a result, in such contexts, meta-evaluation remains a significant hurdle, demanding creative solutions that compensate for the lack of standardized resources.
@@
In the absence of standardized benchmarks and expert annotations, meta-evaluation in domain-specific contexts often relies on creative, task-specific strategies. One illustrative approach is to use outputs of known or controlled quality, such as responses from models with varying capabilities, and verify that the LaaJ assigns higher scores to the better outputs. However, due to the specificity of the domain, there is often little to no literature guiding which metrics are effective or which experimental setups yield meaningful insights for meta-evaluation [hada2024metal]. Even when a metric appears to correlate with human judgment (though such correlation is rarely verifiable), determining a threshold that signifies sufficient evaluator quality remains inherently difficult. Without clear standards, practitioners are left to define ad hoc criteria, which may vary widely across applications and compromise reproducibility.
@@
To address this challenge, we propose LaaJMeter, a simulation-based framework for controlled meta-evaluation of LaaJs. LaaJMeter generates synthetic data representing virtual models and virtual LaaJs, enabling systematic analysis of evaluation metrics under realistic conditions. By simulating models of varying quality and LaaJs with different noise and bias profiles, we can test the sensitivity and robustness of metrics without relying on human annotation [lu2025prompting]. In Section III, we illustrate the problem we address. In Section IV, we describe the LaaJMeter framework.
We demonstrate the utility of LaaJMeter through a use case in code translation, where an LLM translates legacy code into a modern programming language. The IBM watsonx Code Assistant for Z [WCA4Z] (WCA4Z) is IBM’s code assistant for software development on mainframe platforms. A distinctive feature of WCA4Z is its support for mainframe programming languages, such as the legacy language COBOL [COBOL]. A recently developed IBM service transforms COBOL code into the modern Java language [C2J].
@@
As part of the evaluation of the COBOL translation to Java service, we developed several LaaJs and used the LaaJMeter framework to calibrate meta-evaluation of the LaaJs. Specifically, we simulate the metrics: t-test, Kendall’s $\tau$ correlation, and an ordering experiment. Our results reveal that while some metrics (t-test) are insufficiently sensitive to LaaJ quality, others ($\tau$-correlation) provide robust signals even when model distance estimates are imprecise. The ordering experiment, though effective, requires careful interpretation of model distance to ensure meaningful results. We remark that we do not claim for the same metric efficiency in other domains. This is the core advantage of LaaJMeter: it allows to simulate metric simulation tailored to the specific domain, and thus to enhance confidence of the results. In section V we describe the construction of the synthetic data for the use case. In Section VI, we present the LaaJMeter simulations for the use case and consequences. In Section VII, we discuss distance estimations, and provide several suggestions for such estimations.
@@
In summary, LaaJMeter offers a principled and extensible framework for evaluating LLMs as judges, particularly in settings where annotated data is scarce or unreliable. It complements existing evaluation paradigms and provides practical guidance for metric selection and thresholding. Our work contributes to the growing body of research on LLM-based evaluation and opens new directions for scalable, data-efficient assessment of model judgment quality.
II Related Work
Recent studies increasingly employ synthetic or semi-synthetic data to probe the reliability of LLM-as-a-Judge evaluators. Controlled simulation has emerged as a promising approach: researchers generate model responses with known or parametrically defined correctness, then assess whether a LaaJ reproduces the expected quality [kour2022measuringmeasuringtoolsautomatic, fandina2025automatedvalidationllmbasedevaluators]. This approach appears in several recent frameworks for evaluating factuality, reasoning, and preference alignment, where the ground truth is either human-defined or synthetically constructed [tan2025judgebenchbenchmarkevaluatingllmbased, fandina2025safesafetymetricautomatic]. Such works validate judges themselves, that is, they test whether a given LaaJ behaves as a reliable evaluator under known or controlled conditions.
When multiple judges are available, prior studies typically rely on conventional statistical decision metrics to select the best judge or to assess relative quality. Common approaches include mean-difference significance tests such as the t-test or ANOVA, correlation-based comparisons with baseline scores, inter-rater agreement statistics, and correlation with human judgments [sai2020surveyevaluationmetricsused, chen2024humansllmsjudgestudy, dong2024llmpersonalizedjudge]. However, human-annotated data are often scarce or domain-specific, motivating the need for tailored techniques [fandina2025vintagecodemodernjudges] or even human-free methods to compare and select evaluators [chern2024l].
To our knowledge no prior work has validated the decision metrics themselves, namely, tested whether the statistical rules used to select among judges reliably identify the better evaluator under known ground-truth conditions. Our framework provides the first controlled, label-free setting for such analysis. demonstrating, for example, that a simple t-test can yield significant results even for poor judges when model differences are large, and thus cannot serve as a reliable criterion for judge selection.
III Problem Description
@@
In the absence of a comprehensive expert evaluation, one way to assess LaaJ quality is by applying it to outputs of known, varying quality. The evaluated LaaJ is expected to correctly distinguish between different outputs to the same input, by assigning higher scores to better outputs. This form of meta-evaluation enables comparison between different LaaJs based on their ability to rank outputs in accordance with their true quality. See [fandina2025automatedvalidationllmbasedevaluators] for an example of such an evaluation. However, even when relative comparisons are possible, it remains difficult to define a satisfactory threshold for what constitutes acceptable LaaJ quality.
@@
IV LaaJMeter, a General Description
To address the problem described above, we propose the creation of synthetic data representing virtual models and virtual LaaJs. This data is designed to simulate the behavior of models and LaaJs on a specific task, enabling controlled meta-evaluation. The key design principle is to ensure that the generated data closely resembles a realistic distribution and the behaviors observed in the evaluated NLP task. In Section V, we present a use case that demonstrates the general approach outlined here. This construction example may help clarify the abstract concepts introduced in this section, which readers might otherwise find difficult to grasp.
IV-A Virtual Points
@@ IV-B Virtual Models
IV-C Additional Virtual Models
To simulate models of varying quality, we generate additional virtual models by systematically modifying the scores of an existing model. For example, given a model $M$, we can define a new model $M^{\prime}$ such that $M^{\prime}(k)=M(k)+a_{k}$, where the score adjustment $a_{k}$ is sampled from a distribution. If the distribution tends to be positive, $M^{\prime}$ represents a model of higher quality than $M$. Importantly, this approach maintains a pointwise relationship between models, ensuring that $M^{\prime}$ performs better than $M$ in expectation across all points (assuming $a_{k}$ has a positive expected value), mimicking the behavior of improved model versions.
IV-D Virtual LaaJs
@@
Table I: $p$-value results
V Code Translation Use Case
The IBM watsonx Code Assistant for Z (WCA4Z) is IBM’s code assistant for software development on mainframe platforms. WCA4Z supports mainframe programming languages, such as the legacy language COBOL. A recently developed IBM service transforms COBOL code into the modern Java language (abbreviated as C2J). As part of the evaluation of the C2J service, we developed several LaaJs and used the LaaJMeter framework to assess their quality. In this section, we present our simulations and findings, thereby demonstrating an embodiment of the abstract concepts outlined in the previous section.
V-A Virtual Models
We began by constructing a virtual base model $M_{0}:\{0,\dots,99\}\rightarrow[0,30]$. This model was created by assigning ground-truth scores to point indices, sampled from the score distribution of a LaaJ we developed for the code translation evaluation. Accordingly, we chose the range $[0,30]$ to match the scoring scale used by that LaaJ. The number of points, 100, is typical to a benchmarks we use for model evaluation.
Following the construction of $M_{0}$, we generated additional virtual models $M_{1},\dots,M_{20}$. Each model $M_{i+1}$ was derived from $M_{i}$ by modifying its scores as follows: for each point $k$, with probability $p$, we set $M_{i+1}(k)=\min\{M_{i}(k)+1,30\}$, incrementing the score unless it exceeds the maximum value of 30. With probability 1p1-p, we set $M_{i+1}(k)=\max\{M_{i}(k)-1,0\}$, decrementing the score unless it falls below 0. The value of $p$ was chosen such that the expected mean score of $M_{i+1}$ is 0.5 points higher than that of $M_{i}$, accounting for the number of scores at the boundaries (0 and 30).
V-B Virtual LaaJs
We now describe the construction of virtual LaaJs, informed by our experience evaluating LaaJs in the code modernization task. In practice, benchmarks often contain a mix of straightforward and challenging examples. A competent LaaJ typically performs well on simple cases, while its performance may vary on examples involving specific features, such as uncommon commands, that introduce evaluation difficulty. Some LaaJs may handle certain features effectively but struggle with others. Practical examples to such special features are the use of CICS commands, IMS commands, SQL queries etc.
VI Code Translation Use Case: Metric Simulation and Findings
VI-A $t$-Test